OpenStudy (anonymous):

Water flows through a rubber hose 2 cm in diameter at a velocity of 2 m/s. What is the rate of flow in cubic meters per seconds?

OpenStudy (anonymous):

The textbook's explanation says pid^2/4 for some reason. I don't have the slightest idea where 4 came from

OpenStudy (anonymous):

Perhaps that is the official formula to get the area of the circle instead of from the radius.

OpenStudy (anonymous):

I'd never been accustomed to using this route instead of radius so.

OpenStudy (anonymous):

No, the volume of a circle is pi times radius squared. You can work out the area of the circle using this if you halve the diameter to get the radius.

OpenStudy (anonymous):

I guess this sums up to (1cm)^2*pi=3.14cm^2=3.14m^2*10^-3*2m/s=6.28*10^-3m^3/s=rate of flow.

OpenStudy (anonymous):

You mean the area of the circle

OpenStudy (anonymous):

Thanks anyway